Rhonda Brennan found her first job after graduating from college through the classifieds of the Miami Herald. She was delighted when the offer came through at $15.30 per hour. She completed her W-4 stating that she is married with a child and claims an allowance of 3. Her company will pay her biweekly for 80 hours (assume a tax rate of 6.2% on $118,500 for Social Security and 1.45% for Medicare). Calculate her take-home pay for her first check. (Use Table 9.1 and Table 9.2) (Round your answer to the nearest cent.)
Answer #1

No of hours worked= 80

Salary per hour= $15.30

A.Therefore, total salary she will be receiving for 80 hours is 80*15.30 = $1224

Now total deductions is 6.2% and 1.45%

B.So total deductions is $93.63

Total take home pay for her first check= A-B= 1224-93.63 = $1184.37